OpenStudy (anonymous):

If sinA=5/6, where angle A is in Quadrant 1, find the exact value of sin (1/2A)

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

you're asked to find \[\sin( \frac{1}{2}A)\]?or,\[\sin( \frac{1}{2A})\]?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

sin (1/2 A)

OpenStudy (anonymous):

you can use the half angle identity\[\sin \frac{A}{2}=\sqrt{\frac{1-\cos A}{2}}\]where,\[\cos A=\sqrt{1-\sin ^2A}\]so\[\cos A=\sqrt{1-\frac{25}{36}}=\frac{\sqrt{11}}{6}\]plug that value of cosine A into the half angle formula to get your answer :)

OpenStudy (anonymous):

What's the final answer?